OpenStudy (anonymous):

Help! I am breaking down an assignment. So can you help with this part anyone? Andrew makes $6 an hour plus $9 an hour for every hour of overtime. Overtime hours are any hours more than 40 hours for the week. Part A: Create an equation that shows the amount of wages earned, W, for working x hours in a week when there is no overtime.

OpenStudy (anonymous):

i would help if it was "HelpPrettyChic" ;)

OpenStudy (anonymous):

hmm

OpenStudy (countryboy15):

we just did this

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

did you not understand it?

OpenStudy (countryboy15):

A) w(t) = 6 t { t< or = 40 } B) if t is greater than 40, Andrew makes 240 for is full time pay + $9 for every hour over 40 w(t) = 6 t { t< or = 40 } and = 240 + 9 (t-40) { t>40 you can simplify the second equation to w(t) = -120 + 9t {t >40 C) Andrew made 330. 330>240 so t> 40 -120 + 9t = 330 9t = 450 t = 50

OpenStudy (countryboy15):

hope this helps

OpenStudy (anonymous):

exacty

OpenStudy (countryboy15):

lol

OpenStudy (anonymous):

medals ll around

OpenStudy (anonymous):

Where is the 'T' commin from?

OpenStudy (anonymous):

I don't understand why there is a 't' in your equations???

OpenStudy (anonymous):

did you get the questions correct?

OpenStudy (anonymous):

t was just a simple variable to remeber what we were doing

OpenStudy (anonymous):

Well its written out response, I did not submit yet..

OpenStudy (anonymous):

jdi

OpenStudy (anonymous):

JUST DO IT

OpenStudy (anonymous):

I want to see if you could write a equation according to what I out with no extra variable though. i got in trouble last time I did that ):

OpenStudy (anonymous):

Can anybody do that please?? Just write the equations?

OpenStudy (anonymous):

I was right! Those answers were not right.. I found the right one!

OpenStudy (anonymous):

Any other way you'd write it?